4. A constant force of 45N acts on a 15 kg box. The direction of the force is to the left. If the box is originally at rest, what will be its velocity after 5s? Indicate the magnitude and direction of the velocity.

1
Expert's answer
2022-01-23T16:17:13-0500

Given:

"F=45\:\rm N"

"m=15\:\rm kg"

"t=5\:\rm s"

"v_0=0\:\rm m/s"


The acceleration of a box

"a=\frac{F}{m}=\frac{45}{15}=3\:\rm m/s^2"

The velocity of a box after "5\:\rm s"

"v=v_0+at=0+3*5=12\:\rm m/s\quad (toward\; left)"